OpenStudy (anonymous):

8. One more rectangular-shaped piece of metal siding needs to be cut to cover the exterior of a pole barn. The area of the piece is 30 ft ^2. The length is 1 less than 3 times the width. How wide should the metal piece be? Round to the nearest hundredth. A. 3.33 ft B. 4.3 ft C.1ft D.30 ft

OpenStudy (anonymous):

@rishavraj

OpenStudy (anonymous):

@rational

OpenStudy (anonymous):

@ananthus123

OpenStudy (anonymous):

@texaschic101

OpenStudy (texaschic101):

Area of a rectangle : A = L * W A = 30 L = 3W - 1 (length is 1 less then 3 times width) now we sub 30 = (3W - 1)(W) 30 = 3W^2 - W 3w^2 - W - 30 = 0 quadratic formula : x = (-b (+-) sq rt b^2 - 4(a)(c)) / (2a) a = 3, b = -1, c = -30 x = -(-1) (+-) sq rt -1^2 - 4(3)(-30) / (2(3) x = 1 (+-) sq rt 1 + 360) / 6 x = 1 (+-) sq rt 361/6 x = 1 (+-) 19/6 x = (1 + 19)/6 = 20/6 = 3 1/3 = 3.333 <== your answer x = (1 - 19) / 6 = -18/6 = -3 <== extraneous solution

OpenStudy (anonymous):

Thx @texaschic101

OpenStudy (texaschic101):

not a problem...happy to help :)